# Friend or Foe? Harnessing Controllable Overfitting for Anomaly Detection

### v1

@@ 3.3.2 Introduction of RADI
where $S_{a}$ and $S_{n}$ represent the scores of anomalous and normal pixels, respectively.
It might seem counterintuitive why using RADI as an indicator would enhance model performance, as one might think that maximizing this value would lead to the model predicting everything as anomalous. However, RADI itself is not directly used for model optimization; rather, it serves as a metric to reflect the optimization outcome and help identify the extent to which normal and anomalous distributions are distinguishable. The goal is to ensure that anomalies receive consistently higher scores compared to normal samples, which allows the model to achieve better discrimination.
Moreover, the relationship between overfitting and variance is controlled by  Eq. 8, which ensures that the variance of normal pixel predictions ($\sigma_{n}(\theta)$) does not decrease to zero due to the noise term $\sigma_{\text{noise}}(\theta)$. This prevents the variance from collapsing entirely, maintaining a distinction between normal and anomalous scores. Thus, RADI can increase without the model predicting all samples as anomalous, as variability in normal predictions remains above a non-zero threshold, sustaining a balance between sensitivity to anomalies and generalization.
@@
3.3.3 Theoretical Derivation of Optimal Overfitting Conditions
For our case, where both $S_{a}$ and $S_{n}$ follow normal distributions, RADI can be further expressed in a simplified analytical form:
$\textit{RADI}(\theta)=\Phi\left(\frac{\mu_{a}-\mu_{n}}{\sqrt{\sigma_{n}(\theta)^{2}+\sigma_{a}^{2}}}\right),$ (7)
where $\Phi$ is the CDF of the standard normal distribution. In this case, $\mu_{a}$ and $\mu_{n}$ are the means of the anomaly and normal pixel score distributions, respectively, while $\sigma_{a}$ and $\sigma_{n}(\theta)$ are their respective standard deviations, with $\theta$ representing the Aberrance Retention Quotient.
@@ 3.3.3 Theoretical Derivation of Optimal Overfitting Conditions
$\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta}+\sigma_{\text{noise}}(\theta),$ (8)
@@
$\sigma_{\text{noise}}(\theta)=\sigma_{\text{max}}(1-e^{-h(\theta-\theta_{0})}),\quad\text{for}\;\theta>\theta_{0},$ (9)
@@
Category DSR [31] PatchCore [21] BGAD [26] SimpleNet [17] GLASS [4] RD [6] RD (+COAD) RD++ [23] RD++ (+COAD)
Object Bottle 99.6/98.8 100/98.5 100/98.9 100/98.0 100/99.3 100/98.7 100/98.8 100/98.8 100/99.6
@@
Table 1: Performance comparison across different methods in one-class tasks. Bold text indicates the best performance among all methods, while underlined text shows the comparison between our improved method and the original method. The values in the form of xx/xx represent image-level AUROC / pixel-level AUROC.
@@
$\frac{d\textit{RADI}(\theta)}{d\theta}=\phi(z(\theta))\cdot\frac{dz(\theta)}{d\theta}=0,$ (10)
@@
$\frac{dz(\theta)}{d\theta}=-\frac{(\mu_{a}-\mu_{n})\cdot\sigma_{n}(\theta)\cdot\frac{d\sigma_{n}(\theta)}{d\theta}}{(\sigma_{n}(\theta)^{2}+\sigma_{a}^{2})^{3/2}}=0,$ (11)
@@
$\theta^{*}=\frac{\ln(k\sigma_{n0})-\ln(h\sigma_{\text{max}})+h\theta_{0}}{k-h}.$ (12)
3.3.4 Dual Control Mechanism and Freeze Command Signals
We define Aberrance Retention Quotient of $\textit{ARQ}=\theta$, adjusted to optimize the pixel-level AUROC. The mathematical relationship between ARQ and AUROC is established by monitoring the gradient, We call it Gradient-Guided Overfitting Control.
$\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0.$ (13)
@@ 3.3.4 Dual Control Mechanism and Freeze Command Signals
$\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0,\quad\text{for}\;\theta\in\textit{$ARQ_{optimal}$}.$ (14)
If Dual Control Mechanism fails, the Dual Control Mechanism triggers Freeze Command Signals, which prompt whether freeze certain network layers, beginning with lower-level feature extraction layers and moving upwards.
The rationale for progressively freezing certain layers while overfitting the other parts lies in the architecture’s functional partitioning. Lower layers are responsible for extracting fundamental features, which are common across both normal and anomalous data. Freezing these layers early preserves their ability to extract generalizable features, preventing degradation due to overfitting. Conversely, keeping deeper layers unfrozen during the overfitting stage enables the model to specialize in identifying unique, high-level features that are more discriminative between normal and anomalous samples. This approach helps maintain generalization while enhancing sensitivity to rare patterns, ultimately boosting anomaly detection performance without succumbing to the pitfalls of overfitting the entire model.
4 Experiments
Category US [2] PSVDD [27] PaDiM [5] MKD [22] DRAEM [30] UniAD [28] UniAD (+COAD) DiAD [11] DiAD (+COAD)
Object Bottle 84.0/67.9 85.5/86.7 97.9/96.1 98.7/91.8 97.5/87.6 99.7/98.1 100/98.0 99.7/98.4 100/98.8
@@ 4 Experiments
Table 2: Performance comparison across different methods in multi-class tasks. Bold text indicates the best performance among all methods, while underlined text shows the comparison between our improved method and the original method. The values in the form of xx/xx represent image-level AUROC / pixel-level AUROC.
4.1 Dataset
We evaluate our method using the MVTec AD dataset [1], which is a comprehensive, real-world dataset for unsupervised anomaly detection tasks. The MVTec AD dataset includes 15 categories, comprising 3629 images for training and 1725 images for testing. The dataset provides pixel-level annotations for each anomaly, making it suitable for evaluating both image-level classification and pixel-level segmentation tasks.
4.2 Comparison with SOTAs
In this section, we compare the performance of our proposed method with several SOTA anomaly detection methods in both one-class and multi-class tasks.
Using the optimal ARQ value $\theta$ derived from the $ARQ_{o}ptimal$ as indicated by  Eq. 12, we set ARQ to 0.006 for one-class tasks and 0.06 for multi-class tasks. $ARQ_{optimal}$ is defined as $\theta\in[0.001,0.011]$ for one-class tasks and $\theta\in[0.01,0.11]$ for multi-class tasks, where the balance between overfitting and generalization is optimal for anomaly detection. We saved model states within these intervals and evaluated their performance across respective benchmarks to ensure robust detection.
Furthermore, during model training, we adopted a Dual Control Mechanism for Optimal Detection Performance approach in our framework COAD, as detailed in  Eq. 14, and when Dual Control Mechanism fails, it triggers Freeze Command Signals, prompting the selective freezing of certain network layers in the traditional framework.
@@ 4.2 Comparison with SOTAs
The results of our one-class and multi-class experiments are summarized in  Tabs. 1 and 2. It can be observed that our enhanced method outperforms the original methods, achieving new SOTA results. The enhancement in pixel-level AUROC is especially significant, which is crucial for fine-grained anomaly detection.
These results confirm that the introduction of controllable overfitting, regulated by ARQ and RADI, enhances the model’s sensitivity and stability, leading to robust anomaly detection abilities that outperform existing SOTA methods.
4.3 Ablation Study
4.3.1 Comparisons with Our Basic Frameworks
@@
Table 3: Image-Level AUROC comparison between the basic and enhanced versions of different frameworks.
@@
Table 4: Pixel-Level AUROC comparison between the basic and enhanced versions of different frameworks.
We conduct an ablation study comparing our enhanced framework with its corresponding basic versions in  Tabs. 3 and 4. The results demonstrate consistent and notable improvements in both image-level and pixel-level AUROC across all evaluated frameworks.
In particular, the performance gains observed in image-level anomaly detection emphasize the strength of our COAD, which effectively captures subtle differences between normal and anomalous regions. Specifically, our enhancements yielded an increase of up to 1.9 in the DiAD framework, as seen in image-level AUROC, and similar significant improvements in other frameworks.
For pixel-level AUROC, a similar pattern of improvement was observed, with gains of up to 1.2 seen in multiple frameworks, as shown in Tab. 4. This underscores the strength of our COAD, which not only optimizes image-level detection but also provides enhanced granularity in distinguishing pixel-level abnormalities.
Furthermore, the versatility of our method is evident in its applicability to different anomaly detection paradigms. Whether in one-class or multi-class anomaly detection tasks, or even in settings employing diffusion-based models versus non-diffusion frameworks, our method has consistently demonstrated superior performance, showcasing our method’s robustness and generalizability.
4.3.2 Comparisons with Distinct ARQ
@@
Table 5: Comparison of performance across distinct ARQ ranges for different models in one-class tasks.
@@
Table 6: Comparison of performance across distinct ARQ ranges for different models in multi-class tasks.
In this section, we compare the performance of different models under distinct ARQ ranges to validate the impact of controllable overfitting on anomaly detection. We evaluated methods such as RD, RD++, UniAD, and DiAD across various ARQ levels. As  Tabs. 5 and 6 show, the results demonstrate that, by appropriately regulating the ARQ, which is in $ARQ_{optimal}$, model detection performance can be significantly enhanced. It is evident that our controllable overfitting strategy not only yields consistent improvements across different models but also helps in striking the optimal balance between overfitting and generalization.
@@
4.4 Validation of Distribution Assumptions and Theoretical Validation of Gaussian Noise Usage
@@
Table 7: Partial Results of Validation Metrics for Assumptions on Normal and Anomalous Pixel Score Distributions.
In this section, we firstly validate the assumptions introduced in Sec. 3.3.1 by employing the Total Variation Distance (TVD) metric. We leveraging multiple models at different stages of overfitting, characterized by distinct ARQ values. For each model state, corresponding to different ARQ = $\theta$, we calculate the prediction scores for both normal and anomalous pixels across the dataset. Subsequently, we evaluate their proximity to a Gaussian distribution using TVD, ensuring a robust assessment of the normality of these distributions. The partial results of this validation are summarized in Tab. 7. Moreover, we model the variance of the normal pixel prediction scores $(\sigma_{n}(\theta))$ as an exponential function of the ARQ = $\theta$, given by Eq. 4.
To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we employ the TVD metric to quantitatively compare the distribution of real anomaly scores with that of Gaussian-distributed pseudo-anomalies. We can see  Tab. 7, the resulting TVD value was about 0.08, suggesting that the Gaussian-distributed pseudo-anomalies exhibit a significant overlap with the actual anomalies. This supports the assumptions that Gaussian noise can be effectively used to preliminarily simulate anomalous behavior in training models for anomaly detection, which is usually used by GLASS[4], SimpleNet[17], RealNet[32],and DDPM [12].
5 Conclusion
In this work, we reenvisioned overfitting as a controllable and transformative mechanism, using COAD to enhance model capabilities beyond conventional boundaries. We introduced ARQ to precisely regulate overfitting, as well as RADI, which leverages $ARQ_{optimal}$ to provide a more versatile metric compared to AUROC-pixel, facilitating both theoretical modeling and Dual Control Mechanism. By repurposing overfitting as a generalizable module, our approach not only dismantles its demonization but also enables us to achieve SOTA results in both one-class and multi-class anomaly detection tasks. Furthermore, we provide a robust theoretical foundation for employing Gaussian noise as a preliminary pseudo-anomaly generator, extending the applicability of our COAD.
Supplementary Material
6 Dual Control Mechanism Design Concept

### v2

@@ 3.3.2. Introduction of RADI
where $S_{a}$ and $S_{n}$ represent the scores of anomalous and normal pixels, respectively. Intuitively, RADI quantifies the probability that a selected anomalous pixel has a higher anomaly score than a selected normal pixel. A higher RADI thus indicates clearer model discrimination between anomalies and normal patterns.
Although maximizing RADI might seem to encourage over-classifying anomalies, in practice, the variance of normal pixel scores remains above a non-zero lower bound due to controlled noise ( eq. 8). This ensures normal samples retain distinct scores from anomalies, preventing the trivial solution of classifying all samples as anomalous.
@@
3.3.3. Theoretical Derivation of Optimal Overfitting Conditions
For our case, where both $S_{a}$ and $S_{n}$ follow normal distributions, to analytically derive the optimal level of controllable overfitting, we explicitly express RADI in closed-form under normal distribution assumptions. This facilitates understanding how statistical changes induced by overfitting influence anomaly detection performance:
(7) $\textit{RADI}(\theta)=\Phi\left(\frac{\mu_{a}-\mu_{n}}{\sqrt{\sigma_{n}(\theta)^{2}+\sigma_{a}^{2}}}\right),$
where $\Phi$ is the CDF of the standard normal distribution. In this case, $\mu_{a}$ and $\mu_{n}$ are the means of the anomaly and normal pixel score distributions, respectively, while $\sigma_{a}$ and $\sigma_{n}(\theta)$ are their respective standard deviations, with $\theta$ representing the ARQ.
@@ 3.3.3. Theoretical Derivation of Optimal Overfitting Conditions
(8) $\sigma_{n}(\theta)=\sigma_{n0}e^{-k\theta}+\sigma_{\text{noise}}(\theta),$
@@
(9) $\sigma_{\text{noise}}(\theta)=\sigma_{\text{max}}(1-e^{-h(\theta-\theta_{0})}),\quad\text{for}\;\theta>\theta_{0},$
@@
Table 1. Performance comparison across different methods in one-class tasks on MVTec AD dataset. Bold text indicates the best performance among all methods, while underlined text shows the comparison Method vs. Method (+COAD). The values in the form of xx/xx represent image-level AUROC / pixel-level AUROC.
Category DSR (Zavrtanik et al., 2022) PatchCore (Roth et al., 2022) BGAD (Yao et al., 2023) SimpleNet (Liu et al., 2023) GLASS (Chen et al., 2024) RD (Deng and Li, 2022) RD (+COAD) RD++ (Tien et al., 2023) RD++ (+COAD)
Object
Bottle 99.6/98.8 100/98.5 100/98.9 100/98.0 100/99.3 100/98.7 100/98.8 100/98.8 100/99.6
@@
Table 2. Performance comparison across different methods in one-class tasks on VisA dataset.
Category PatchCore (Roth et al., 2022) SimpleNet (Liu et al., 2023) DRAEM (Zavrtanik et al., 2021) RD (Deng and Li, 2022) RD(+COAD) RD++ (Tien et al., 2023) RD++(+COAD)
Single Instance Cashew 97.7/99.2 94.8/99.0 88.3/85.0 99.6/95.8 97.8/97.6 97.8/95.8 98.1/98.3
Chewinggum 99.1/98.9 100/98.5 96.4/97.7 99.7/99.0 97.8/99.4 96.4/99.4 97.4/99.5
Fryum 91.6/95.9 96.6/94.5 94.7/82.5 96.6/94.3 95.5/96.3 95.8/96.5 96.8/98.1
Pipe fryum 99.0/99.3 99.2/99.3 94.7/65.8 97.0/92.0 99.4/99.1 99.6/99.1 99.3/99.4
Multiple Instances Candle 98.7/99.2 96.9/98.6 89.6/91.0 92.2/97.9 95.1/99.0 96.4/98.6 92.8/98.7
Capsules 68.8/96.5 89.5/99.2 89.2/99.0 90.1/89.5 87.5/99.3 92.1/99.4 92.7/99.7
Macaroni1 90.1/98.5 97.6/99.6 93.9/99.4 98.4/97.7 95.1/99.3 94.0/99.7 95.7/99.9
Macaroni2 63.4/93.5 83.4/96.4 88.3/99.7 97.6/87.7 96.6/99.3 88.0/99.7 92.5/99.6
Complex Structure PCB1 96.0/99.8 99.2/99.8 84.7/98.4 97.6/75.0 96.5/99.7 97.0/99.7 98.8/99.5
PCB2 95.1/98.4 99.2/98.8 96.2/94.0 91.1/64.8 94.0/98.8 97.2/99.0 97.3/98.9
PCB3 93.0/98.9 98.6/99.2 97.4/94.3 95.5/95.5 95.9/99.3 96.8/99.2 97.4/99.3
PCB4 99.5/98.3 98.9/98.6 98.8/97.6 96.5/92.8 99.9/98.7 99.8/98.6 98.5/99.6
Avg. 91.0/98.1 96.2/98.5 92.4/92.0 96.0/90.1 95.9/98.8 95.9/98.7 96.4/99.2
@@
(10) $\frac{d\textit{RADI}(\theta)}{d\theta}=\phi(z(\theta))\cdot\frac{dz(\theta)}{d\theta}=0,$
@@
(11) $\frac{dz(\theta)}{d\theta}=-\frac{(\mu_{a}-\mu_{n})\cdot\sigma_{n}(\theta)\cdot\frac{d\sigma_{n}(\theta)}{d\theta}}{(\sigma_{n}(\theta)^{2}+\sigma_{a}^{2})^{3/2}}=0,$
@@
(12) $\theta^{*}=\frac{\ln(k\sigma_{n0})-\ln(h\sigma_{\text{max}})+h\theta_{0}}{k-h}.$
3.3.4. Dual Control Mechanism and Freeze Command Signals
We define Aberrance Retention Quotient of $\textit{ARQ}=\theta$, adjusted to optimize the pixel-level AUROC. The mathematical relationship between ARQ and AUROC is established by monitoring the gradient, we call it Gradient-Guided Overfitting Control. The Dual Control Mechanism leverages both ARQ and RADI simultaneously: ARQ ensures overfitting is maintained within beneficial bounds, while RADI’s gradient condition ensures the continual improvement of anomaly detection performance.
(13) $\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0.$
@@ 3.3.4. Dual Control Mechanism and Freeze Command Signals
(14) $\frac{d\textit{RADI}(\theta)}{d\theta}\geq 0,\quad\text{for}\;\theta\in\textit{$ARQ_{optimal}$}.$
When ARQ exceeds its optimal bounds and RADI gradient condition fails, Freeze Command Signals are activated, progressively freezing layers starting from the earliest layers upwards.
The rationale for progressively freezing certain layers while overfitting the other parts lies in the architecture’s functional partitioning. Lower layers are responsible for extracting fundamental features, which are common across both normal and anomalous data. Freezing these layers early preserves their ability to extract generalizable features, preventing degradation due to overfitting. Conversely, keeping deeper layers unfrozen during the overfitting stage enables the model to specialize in identifying unique, high-level features that are more discriminative between normal and anomalous samples. This approach ultimately boosts performance without succumbing to the pitfalls of overfitting the entire model.
4. Experiments
Table 3. Performance comparison across different methods in multi-class tasks on MVTec AD dataset.
Category US (Bergmann et al., 2020) PSVDD (Yi and Yoon, 2020) PaDiM (Defard et al., 2021) MKD (Salehi et al., 2021) DRAEM (Zavrtanik et al., 2021) UniAD (You et al., 2022) UniAD (+COAD) DiAD (He et al., 2023) DiAD (+COAD)
Object
Bottle 84.0/67.9 85.5/86.7 97.9/96.1 98.7/91.8 97.5/87.6 99.7/98.1 100/98.0 99.7/98.4 100/98.8
@@ 4. Experiments
We evaluate our method using the MVTec AD dataset (Bergmann et al., 2019) and VisA dataset (Zou et al., 2022), which are comprehensive, real-world datasets for unsupervised anomaly detection tasks. The MVTec AD dataset includes 15 categories, comprising 3629 images for training and 1725 images for testing. The VisA dataset includes 12 categories, including 9,621 normal and 1,200 anomalous samples.
4.1. Comparison with SOTAs
In this section, we compare the performance of our proposed method with several SOTA anomaly detection methods in both one-class and multi-class tasks on MVTec AD and VisA datasets.
Using the optimal ARQ value $\theta$ derived from the $ARQ_{optimal}$ as indicated by  eq. 12, we set ARQ to 0.006 for one-class tasks and 0.06 for multi-class tasks as shown in  section 4.2.2 which is chosen empirically. $ARQ_{optimal}$ is defined as $\theta\in[0.001,0.011]$ for one-class tasks and $\theta\in[0.01,0.11]$ for multi-class tasks, where the balance between overfitting and generalization is optimal for anomaly detection.
Furthermore, during model training, we adopt a Dual Control Mechanism for Optimal Performance method in our framework COAD, as detailed in  eq. 14, and when Dual Control Mechanism fails, it triggers Freeze Command Signals, prompting selective freezing of certain network layers in the traditional framework.
@@ 4.1. Comparison with SOTAs
The quantitative results of our experiments across both one-class and multi-class anomaly detection tasks are comprehensively summarized in  tables 1, 2 and 3, with additional detailed results for the VisA multi-class task provided in  table 9 of the Supplementary Materials  appendix D. Clearly, our proposed COAD framework consistently surpasses previous SOTA approaches, delivering substantial improvements in both image-level and pixel-level AUROC metrics across all evaluated datasets and tasks. Additionally, partial qualitative visualization results demonstrating the effectiveness of COAD are presented in  figs. 3 and 4, showing high accuracy. Complete qualitative comparisons are provided in  appendix E of the Supplementary Materials.
Figure 3. Qualitative visualization results on partial categories from the MVTec AD for the one-class task. Columns show Original images, GroundTruth, and predictions generated by our COAD framework.
Figure 4. Qualitative visualization results on partial categories from the MVTec AD for the multi-class task.
In conclusion, these comprehensive experimental validations solidify the practical advantages of our COAD framework, highlighting its ability to reliably enhance existing anomaly detection models through controlled and strategic utilization of overfitting.
4.2. Ablation Study
4.2.1. Comparisons with Our Basic Frameworks
Table 4. Image-Level AUROC comparison between the basic and enhanced versions of different frameworks.
@@ 4.2.1. Comparisons with Our Basic Frameworks
VisA Baseline 96.0 95.9 85.5 -
+COAD 95.9(-0.1) 96.4(+0.5) 91.0(+5.5) -
Table 5. Pixel-Level AUROC comparison between the basic and enhanced versions of different frameworks.
@@
VisA Baseline 90.1 98.7 95.9 -
+COAD 98.8(+8.7) 99.2(+0.5) 97.5(+1.6) -
We conduct an ablation study comparing our enhanced framework with its corresponding basic versions in  tables 4 and 5. The results demonstrate consistent and notable improvements in both image-level and pixel-level AUROC across all evaluated frameworks.
In particular, the performance gains observed in image-level anomaly detection emphasize the strength of our COAD, which effectively captures subtle differences between normal and anomalous regions. Specifically, our enhancements yield an increase of up to 1.9 in the DiAD framework, as seen in image-level AUROC, and similar significant improvements in other frameworks.
For pixel-level AUROC, a similar pattern of improvement is observed, with gains of up to 1.2 seen in multiple frameworks, as shown in table 5. This underscores the strength of our COAD, which not only optimizes image-level detection but also provides enhanced granularity in distinguishing pixel-level abnormalities.
Furthermore, the versatility of our method is evident in its applicability to different anomaly detection paradigms. Our method has consistently demonstrated superior performance, showcasing our method’s robustness and generalizability.
4.2.2. Comparisons with Distinct ARQ
Table 6. Comparison of performance across distinct ARQ ranges for different models in one-class tasks on MVTec AD.
@@ 4.2.2. Comparisons with Distinct ARQ
Table 7. Comparison of performance across distinct ARQ ranges for different models in multi-class tasks on MVTec AD.
@@
In this section, we compare the performance of different models under distinct ARQ ranges to validate the impact of controllable overfitting on anomaly detection.
We evaluate methods such as RD, RD++, UniAD, and DiAD across various ARQ levels. As  tables 6 and 7 show, the results demonstrate that, by appropriately regulating the ARQ, which is in $ARQ_{optimal}$, model detection can be significantly enhanced. It is evident that our controllable overfitting strategy not only yields consistent improvements across different models but also helps in striking the balance between overfitting and generalization.
@@
4.3. Validation of Distribution Assumptions and Theoretical Validation of Gaussian Noise Usage
(a) Carpet (Normal)
(b) Wood (Normal)
(c) Carpet (Anomaly)
(d) Wood (Anomaly)
Figure 5. Histograms for normal and anomaly pixel scores of partial categories of MVTec AD dataset. Each histogram includes Gaussian fitting curves for better visualization.
Table 8. Partial Results of Validation Metrics for Assumptions on Normal and Anomalous Pixel Score Distributions.
@@ 4.3. Validation of Distribution Assumptions and Theoretical Validation of Gaussian Noise Usage
In this section, we firstly validate the assumptions introduced in section 3.3.1 by employing the Total Variation Distance (TVD) metric. We leverage multiple models at different stages of overfitting, characterized by distinct ARQ values. For each model state, corresponding to different ARQ = $\theta$, we calculate the prediction scores for all pixels across the dataset. Subsequently, we evaluate their proximity to a Gaussian distribution using TVD. The partial results of this validation are summarized in table 8. Moreover, we model the variance of the normal pixel prediction scores $(\sigma_{n}(\theta))$ as an exponential function of the ARQ, given by eq. 4. Partial qualitative result can be seen in  fig. 5 and complete qualitative study can be seen in  figs. 7 and 6 in Supplementary Materials.
To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we employ the TVD metric to quantitatively compare the distribution of real anomaly scores with that of Gaussian-distributed pseudo-anomalies. We can see  table 8, the resulting TVD value is about 0.08, suggesting that the Gaussian-distributed pseudo-anomalies exhibit a significant overlap with the actual anomalies. This supports the assumptions that Gaussian noise can be used to preliminarily simulate anomalous behavior for anomaly detection, which is usually used by GLASS(Chen et al., 2024), SimpleNet(Liu et al., 2023), RealNet(Zhang et al., 2024),and DDPM (Ho et al., 2020). Extended Discussion of Gaussian Noise Validation can be seen in  appendix C in Supplementary Materials.
5. Conclusion
In this work, we reconceptualize overfitting as a controllable and transformative mechanism, using COAD to enhance model capabilities beyond conventional boundaries. We introduce ARQ to precisely regulate overfitting, as well as RADI, which leverages $ARQ_{optimal}$ to provide a more versatile metric compared to AUROC-pixel, facilitating both theoretical modeling and Dual Control Mechanism. By repurposing overfitting as a generalizable module, our approach not only dismantles its demonization but also enables us to achieve SOTA results in both one-class and multi-class anomaly detection tasks. Furthermore, we provide a robust theoretical foundation for employing Gaussian noise as a preliminary pseudo-anomaly generator, extending the applicability of our COAD.
Appendix
Appendix A Dual Control Mechanism Design Concept